Shift report 20030224 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Mon Feb 24 2003 - 15:49:20 EST

  • Next message: Apache: "Shift report 20030224 08:00-16:00"
    ------------------------------------------------------------
    Date: Mon Feb 24 15:49:20 2003, Shift: 08:00-16:00
    Supervisor: MG, with: 
    Log book#14 pages 35 - 35
    
    Shift Summary:
    ------------------------------------------------------------
    1.Good beam from 09:26. After then background is oscillating because of some studies by MCR,but still between reasonable values.
    2.
    Run 8059: MRS 131 k, FS 45 k 
    Run 8060: MRS 35 k, FS 12 k 
    Run 8061: MRS 192 k, FS 69 k
    Run 8063: no good
    
    TOT: MRS 360 k, FS: 120 k
    
    3. T3 anode A1-vy turned off, because current too high and was tripping all the time. I think it is influenced by C1-yv which also han an unusual high current. In SuperMon T3-1,2 all panels doesn't look as usual, can be because of anode problem. This anode has to be turned off. Because of some bad contacts in slot 18, cumac scaler in crate 2 has been moved to slot 19.D5 is set to error.Called MCR.
    
    PLAN: Continue same settings.
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Mon Feb 24 2003 - 15:53:31 EST